OpenStudy (anonymous):

part 1: Explain, in complete sentences, how you would use the substitution method to solve the following system of equations part 2: provide the solution to the system. x+5y=3 2x+9y=4

OpenStudy (anonymous):

you have to find the value of 1 variable from 1 eq. then substitute that value in other eq. for ex. In, x+5y=3 x=3-5y put this in 2x+9y=4 2(3-5y)+9y=4 Obtain y and the x.